### v1

# Salem numbers and commensurability classes of arithmetic hyperbolic manifolds

Abstract.
In this article we show that given a Salem number $\lambda$, a totally real number field $k\subseteq\mathbb{Q}(\lambda+\lambda^{-1})$, and a positive integer $n\geq\deg_{k}(\lambda)-1$, there exist infinitely many commensurability classes of arithmetic hyperbolic $n$-manifolds defined over $k$ which contain a geodesic of length $\log\lambda$.
The first author was partially supported by NSF grant DMS-2243188
The second author was partially supported by Universal CNPq Grant 408834/2023-4
1. Introduction
A Salem number is a real algebraic integer $\lambda>1$ with the property that all its Galois conjugates except $\lambda^{-1}$ have norm 1. A polynomial $f\in\mathbb{Z}[x]$ is a Salem polynomial if it is the minimal polynomial of a Salem number. Besides the two real roots $\lambda$ and $\lambda^{-1}$ of a Salem polynomial, the remaining roots occur in reciprocal complex-conjugate pairs $e^{i\theta}$ and $e^{-i\theta}$. Salem polynomials are therefore monic of even degree and self-reciprocal, that is $f(x)=x^{\deg(f)}f(x^{-1})$. Although Salem numbers are sometimes defined to have at least one pair of non-real conjugates, we will allow Salem numbers to be quadratic.
Salem numbers appear in many areas of mathematics. One such appearance is related to hyperbolic isometries in arithmetic hyperbolic manifolds of simplest type. If $\ell(\gamma)$ is the translation length of a hyperbolic element $\gamma$ in a classical arithmetic lattice $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$, then $\lambda=e^{\ell(\gamma)}$ is a Salem number. In 2019, Emery, Ratcliffe, and Tschantz established a more direct relationship between Salem numbers and lengths of geodesics in hyperbolic manifolds which are arithmetic of simplest type.
Theorem 1.1.
[ERT19, Theorem 5.2] Let $\Gamma\subseteq\mathrm{Isom}(\mathbb{H}^{n})$ be a classical arithmetic lattice defined over a totally real number field $k$. Let $\gamma$ be a hyperbolic element of $\Gamma$, let $\ell(\gamma)$ be the translation length of $\gamma,$ and let $\lambda=e^{\ell(\gamma)}$. Then
  1. (1)
    $\lambda$ is a Salem number, and $\deg_{k}(\lambda)\leq n+1$
  2. (2)
    $k$ is a subfield of $\mathbb{Q}(\lambda+\lambda^{-1})$, and $[k:\mathbb{Q}(\lambda+\lambda^{-1})]=\deg_{k}(\gamma)/2$.
In the situation of the theorem, we will say that $\Gamma$ realizes $\lambda$. Alternatively, we say that a Salem number $\lambda$ is realized in $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$ if there exists a hyperbolic element $\gamma\in\Gamma$ such that $\lambda=e^{\ell(\gamma)}$.
Emery, Ratcliffe, and Tschantz also showed that every Salem number is realized as follows.
Theorem 1.2.
[ERT19, Theorem 6.3] Let $\lambda$ be a Salem number, let $k$ be a subfield of $\mathbb{Q}(\lambda+\lambda^{-1})$, and let $n\geq 2$ be an integer with $\deg_{k}(\lambda)\leq n+1$. Then there exist a classical arithmetic lattice $\Gamma\subseteq\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ and an orientation-preserving hyperbolic element $\gamma\in\Gamma$ such that $\lambda=e^{\ell(\gamma)}$.
For each Salem number $\lambda$ and number field $k\subseteq\mathbb{Q}(\lambda+\lambda^{-1})$, the arguments in [ERT19] construct a unique such lattice $\Gamma$ in each allowable dimension $n$. This leaves open the following question.
Question.
Let $\lambda$ be a Salem number. Does there exist infinitely many commensurability classes of arithmetic lattices in $\mathrm{Isom}(\mathbb{H}^{n})$ realizing $\lambda$?
The case of $n=2$ and $n=3$ is addressed in [MR03, Corollary 12.2.10]. We answer this question in full generality.
Main Theorem.
Let $\lambda$ be a Salem number, let $k$ be a subfield of $\mathbb{Q}(\lambda+\lambda^{-1})$, and let $n\geq 2$ be an integer with $\deg_{k}(\lambda)\leq n+1$. Then there exist infinitely many commensurability classes of arithmetic lattices in $\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ which realize $\lambda$.
The article is organized as follows. In Section 2 we present the necessary background material on quadratic forms, quaternion algebras, hyperbolic isometries, and arithmetic lattices. This includes a summary of results of Maclachlan on parameterizing commensurability classes of arithmetic lattices of simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$ via quaternion algebras. In Section 3 we gather results of Bayer-Fluckiger giving necessary and sufficient conditions for quadratic spaces to admit an isometry with a given characteristic polynomial. The proof of the Main Theorem comes together in Section 4.
2. Preliminaries
2.1. Quadratic forms over number fields
Let $k$ be a field. A quadratic form over $k$ is a homogeneous polynomial in $r$ variables of degree 2 with coefficients in $k$. It can be written as
$q(x)=\sum_{i=1}^{r}\sum_{j=i}^{r}a_{ij}x_{i}x_{j}=x^{t}S_{q}x,$
where $S_{q}$ is the symmetric matrix
$\begin{pmatrix}a_{11}&\frac{1}{2}a_{12}&\cdots&\frac{1}{2}a_{1r}\\ a_{12}&a_{22}&&\vdots\\ \vdots&&\ddots&\frac{1}{2}a_{r-1,r}\\ \frac{1}{2}a_{1r}&\cdots&\frac{1}{2}a_{r-1,r}&a_{rr}\end{pmatrix}$
with diagonal entries $(S_{q})_{ii}=a_{ii}$ and off-diagonal entries $(S_{q})_{ij}=\frac{1}{2}a_{ij}$. Let $V=k^{r}$, then the quadratic space $(V,q)$ has isometry group given by the orthogonal group
$\mathrm{O}(q)=\mathrm{O}(q)_{k}=\{A\in\mathrm{GL}(r)_{k}\>|\>A^{t}S_{q}A=S_{q}\}.$
The rank of $q$ is the number of variables $\mathrm{rank}(q)=r$. The determinant of $q$ is $\det(q)=\det(S_{q})$. We will assume that $q$ is nondegenerate, that is, $S_{q}$ is invertible, and $\det(q)\not=0$. In what follows, we will only care about quadratic form up to $k^{\times}$-scaling. If $r$ is the rank of $q$, the discriminant of $q$ is $\mathrm{disc}(q)=(-1)^{r(r-1)/2}\det(q)$ as an element of $k^{\times}/k^{\times 2}$, and is thus well-defined up to scaling.
We say that a quadratic form $q$ represents $a\in k$ if there is a vector $v\in K^{n}$ satisfying $q(v)=a$. A quadratic form is isotropic if it represents 0 by a nonzero vector and is anisotropic otherwise.
Two quadratic forms $q$ and $q^{\prime}$ over $k$ are $k$-equivalent if there exists a matrix $M\in\mathrm{GL}(n)_{k}$ such that $M^{t}S_{q}M=S_{q^{\prime}}$. When the field $k$ is clear, we write $q\simeq q^{\prime}$. Any quadratic form over $k$ is $k$-equivalent to a diagonal quadratic form $q^{\prime}$, which means $q^{\prime}(x)=\sum_{i=1}^{n}a_{i}x_{i}^{2}$ and we write $q^{\prime}=\langle a_{1},\dots,a_{n}\rangle$.
Over the reals, any nondegenerate quadratic form is $\mathbb{R}$-equivalent to a diagonal form with coefficients in ±$\pm 1$. The signature of $q$ is the pair $(n_{+},n_{-})$ where $n_{+}$ is the number of $+1$’s and $n_{-}$ is the number of $-1$’s.
Let $k$ be a number field and $\mathcal{O}_{k}$ its ring of integers. We denote by $\Omega$ the set of all places of $k$. The subset of infinite places and finite places will be denoted by $\Omega_{\infty}$ and $\Omega_{fin}$ respectively. The notation “$\nu_{0}$” will denote the identity embedding $k\hookrightarrow\mathbb{R}$. Since the set of finite places of $k$ is in 1-to-1 correspondence with the set of prime ideals in $\mathcal{O}_{k}$, we will abuse notation and often write $\mathfrak{p}$ or Weierstrass-$\wp$ to refer either to prime ideals or to their corresponding finite places.
If $\nu$ is a place of $k$, let $k_{\nu}$ denote the completion corresponding to $\nu$ and let $q_{\nu}$ denote the quadratic form over $k_{\nu}$ obtained from $q$ by $\nu$. If $q$ is defined over a real number field $k\hookrightarrow\mathbb{R}$, we refer to the signature with respect to the identity embedding as the signature of $q$ over $k$.
2.2. Quaternion algebras and invariants of quadratic forms
A quaternion algebra over a field $k$ is a four-dimensional central simple algebra. Any quaternion algebra $B$ can be represented by a Hilbert symbol $(\frac{a,b}{k})$ for some $a,b\in k^{\times}$ where $B$ has a basis $\{1,i,j,ij\}$ with $a,b\in k^{\times}$, $i^{2}=a$, $j^{2}=b$, and $ij=-ji$.
Any quaternion algebra is either isomorphic to the matrix algebra $M_{2}(k)$ or to a division algebra. Over $\mathbb{R}$ or over any local field, there is a unique isomorphism class of division quaternion algebras. Furthermore, $B=(\frac{a,b}{k})\cong M_{2}(k)$ if and only if the quadratic form $ax^{2}+by^{2}-z^{2}$ is isotropic.
Let $k$ be a number field and $\nu$ is a place of $k$. A quaternion algebra $B$ over $k$ is said to be split at $\nu$ if $B_{\nu}\cong M_{2}(k_{\nu})$ and is called ramified at $\nu$ if $B_{\nu}=B\otimes_{k}k_{\nu}$ is a division algebra.
The ramification set of the quaternion algebra $B$ is denoted
$\mathrm{Ram}(B)=\{\nu\in\Omega\>|\>B\text{ is ramified at }\nu\}.$
The subset of infinite ramified places will be denoted by $\mathrm{Ram}_{\infty}(B)$. The subset of finite ramified places will be denoted $\mathrm{Ram}_{fin}(B)$. A Hilbert symbol of a quaternion algebra is not unique. However, quaternion algebras are classified by their ramification sets according to the following theorem (see e.g. [MR03, Theorem 7.3.6]).
Theorem 2.1 (Classification of quaternion algebras).
Let $B$ be a quaternion algebra over a number field $k$. The following hold:
  • $\mathrm{Ram}(B)$ is a finite set of even cardinality.
  • Two quaternion algebras $B_{1}$ and $B_{2}$ over $k$ are isomorphic if and only if $\mathrm{Ram}(B_{1})=\mathrm{Ram}(B_{2})$.
  • If $S$ is any finite subset of $\Omega$ of even cardinality, excluding any complex places, then there exists a quaternion algebra $B$ over $k$ such that $\mathrm{Ram}(B)=S$.
Isomorphism classes of quaternion algebras over $k$ form a subgroup of the Brauer group $\mathrm{Br}(k)$, where every element has order 2. The isomorphism class of the quaternion algebra with empty ramification set corresponds to the identity element in $\mathrm{Br}(k)$. In $\mathrm{Br}(k)$, multiplication of quaternion algebras is given by $B_{1}\cdot B_{2}=B_{3}$ where
$\mathrm{Ram}(B_{3})=\left(\mathrm{Ram}(B_{1})\cup\mathrm{Ram}(B_{2})\right)-\left(\mathrm{Ram}(B_{1})\cap\mathrm{Ram}(B_{2})\right).$
Let $(V,q)$ be a quadratic space over a number field $k$ with diagonalization $\langle a_{1},\dots a_{r}\rangle$. The Hasse invariant of $(V,q)$ is the product
$s(q)=\prod_{i<j}\left(\tfrac{a_{i},a_{j}}{k}\right)\in\mathrm{Br}(k),$
which is independent of the diagonalization, and is invariant in the $k$-equivalence class. This invariant is closely related to the Witt invariant $c(q)$, which can be obtained as follows:
(1) if $\text{if }\mathrm{rank}(q)\equiv 1,2\bmod 8,$ $\>c(q)=s(q)$
if $\text{if }\mathrm{rank}(q)\equiv 3,4\bmod 8,$ $\>c(q)=s(q)\cdot\left(\tfrac{-1,-\det(q)}{k}\right)$
if $\text{if }\mathrm{rank}(q)\equiv 5,6\bmod 8,$ $\>c(q)=s(q)\cdot\left(\tfrac{-1,-1}{k}\right)$
if $\text{if }\mathrm{rank}(q)\equiv 7,8\bmod 8,$ $\>c(q)=s(q)\cdot\left(\tfrac{-1,\det(q)}{k}\right).$
2.3. Hyperbolic space and hyperbolic isometries
The hyperbolic $n$-space $\mathbb{H}^{n}$ is the simply connected Riemannian $n$-manifold in which the sectional curvature is constant $-1$. We will consider the hyperboloid model of hyperbolic space defined as follows.
Let $n\geq 2$ and $q=\langle 1,\dots,1,-1\rangle$ be the standard quadratic form of signature $(n,1)$. The bilinear form $B$ in $\mathbb{R}^{n+1}$ associated with $q$ is given by
$B(x,y)=\frac{q(x+y)-q(x)-q(y)}{2}.$
The set
$\mathcal{S}=\{(x_{1},\ldots,x_{n+1})\in\mathbb{R}^{n+1};q(x_{1},\ldots,x_{n+1})=-1\}$
forms a two-sheeted hyperboloid. We denote the positive sheet by
$\mathcal{S}^{+}=\{(x_{1},\ldots,x_{n+1})\in\mathcal{S};x_{n+1}>0\}.$
The hyperboloid model of the hyperbolic $n$-space is obtained by identifying $\mathbb{H}^{n}$ with the $n$-manifold $\mathcal{S}^{+}$ together with the Riemannian metric given by $d(x,y)=\arccos\left(B(x,y)\right)$.
The orthogonal group $\mathrm{O}(n,1)_{\mathbb{R}}$ preserves the hyperboloid $\mathcal{S}$. We identify the isometry group of $\mathbb{H}^{n}$ with
$\mathrm{Isom}(\mathbb{H}^{n})=\mathrm{PO}(n,1)_{\mathbb{R}}\cong\mathrm{SO}(n,1)_{\mathbb{R}}.$
It has two connected components. The identity component consists of orientation-preserving isometries. The other component consists of orientation-reversing isometries.
We note that one may construct isometric hyperboloid models for $\mathbb{H}^{n}$ using any quadratic form $q^{\prime}$ with signature $(n,1)$ defined over a real field. Since any such $q^{\prime}$ is $\mathbb{R}$-equivalent to the standard form $q=\langle 1,\dots,1,-1\rangle$, there is some $g\in\mathrm{GL}(n+1)_{\mathbb{R}}$ such that $g^{t}S_{q}g=S_{q^{\prime}}$ and $g^{-1}\mathrm{SO}(n,1)_{\mathbb{R}}g=\mathrm{SO}(q)_{\mathbb{R}}$.
An element $\gamma\in\mathrm{Isom}(\mathbb{H}^{n})$ is called hyperbolic if there is a unique geodesic $L\in\mathbb{H}^{n}$, called the axis of $\gamma$, on which $\gamma$ acts as a translation by a positive distance $\ell(\gamma)$, called the translation length of $\gamma$. In this case, $\gamma$ has exactly $n-1$ eigenvalues of norm one, and two real eigenvalues $\lambda=e^{\ell(\gamma)}>1$ and $\lambda^{-1}=e^{-\ell(\gamma)}<1$ not on the unit circle (see [CG74]).
A lattice is a discrete subgroup $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$ such that the quotient $\mathbb{H}^{n}/\Gamma$ has finite volume. Whenever $\Gamma$ is a lattice, the quotient $\mathbb{H}^{n}/\Gamma$ is a complete finite-volume hyperbolic orbifold. In this case, most elements of $\Gamma$ are hyperbolic and the axes of hyperbolic elements map to closed geodesics in the quotient. Also, $\mathbb{H}^{n}/\Gamma$ will be a manifold whenever $\Gamma$ is torsion free.
Two lattices $\Gamma_{1}$ and $\Gamma_{2}$ in $\mathrm{Isom}(\mathbb{H}^{n})$ are commensurable if there exists $g\in\mathrm{Isom}(\mathbb{H}^{n})$ such that $\Gamma_{1}\cap g\Gamma_{2}g^{-2}$ has finite index in both $\Gamma_{1}$ and $g\Gamma_{2}g^{-2}$.
2.4. Arithmetic lattices
Let $k$ be a totally real number field. Consider a quadratic form $q:\mathbb{R}^{n+1}\rightarrow\mathbb{R}$ defined over $k$, with signature $(n,1)$ with respect to the identity embedding $\nu_{0}:k\hookrightarrow\mathbb{R}$.
We say that such a form $q$ is admissible if for any non-identity real place $\nu$ of $k$, $q_{\nu}$ has signature $(n,0)$ (in other words, $q_{\nu}$ is positive definite). This condition on the Galois conjugates of $q$ guarantees that the subgroup
$\mathrm{SO}(q)_{\mathcal{O}_{k}}=\{A\in\mathrm{SL}(n+1)_{\mathcal{O}_{k}};A^{t}S_{q}A=S_{q}\}$
is a discrete subgroup of $\mathrm{SO}(q)_{\mathbb{R}}\cong\mathrm{Isom}(\mathbb{H}^{n})$. Furthermore, $\mathrm{SO}(q)_{\mathcal{O}_{k}}$ is a lattice in $\mathrm{Isom}(\mathbb{H}^{n})$.
Any lattice $\Gamma<\mathrm{Isom}(\mathbb{H}^{n})$ commensurable with some $\mathrm{SO}(q)_{\mathcal{O}_{k}}$ for an admissible $q$ defined over a totally real number field $k$ is called arithmetic of simplest type. The quotient orbifold $\mathbb{H}^{n}/\Gamma$ is also called arithmetic of simplest type.
An arithmetic lattice of simplest type $\Gamma$ is called classical if it is contained in $\mathrm{SO}(q)_{k}$ for an admissible $q$ defined over a totally real number field $k$. It turns out that when $n$ is even, every arithmetic lattice in $\mathrm{Isom}(\mathbb{H}^{n})$ is classical, and necessarily of simplest type (see e.g. [ERT19, Lemma 4.2]). When $n$ is odd, there exist arithmetic lattices that are not of simplest type; however, if $\Gamma$ is arithmetic of simplest type, then $\Gamma^{(2)}$, the subgroup generated by the squares of the elements in $\Gamma$, is classical (see [ERT19, Lemma 4.5]).
In [Mac11], Maclachlan classified the commensurability classes of arithmetic lattices of simplest type in terms of the Witt invariant of quadratic forms. The following statement is a direct consequence of the proofs of [Mac11, Theorems 7.2 and 7.4].
Proposition 2.2.
[Mac11] Let $\Gamma=\mathrm{SO}(q)_{\mathcal{O}_{k}}$ and $\Gamma^{\prime}=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$ be arithmetic lattices of the simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$, where $q$ and $q^{\prime}$ are two admissible quadratic forms defined over a totally number real field $k$.
If $n$ is even, then $\Gamma$ and $\Gamma^{\prime}$ are commensurable if and only if $c(q)=c(q^{\prime})$.
If $n$ is odd, let $\delta=\mathrm{disc}(q)$ and $\delta^{\prime}=\mathrm{disc}(q^{\prime})$. Then $\Gamma$ and $\Gamma^{\prime}$ are commensurable if and only if $\delta=\delta^{\prime}$ and $c(q)\otimes_{k}k(\sqrt{\delta})=c(q^{\prime})\otimes_{k}k(\sqrt{\delta})$ in $\mathrm{Br}(k(\sqrt{\delta}))$.
Let $\{\mathfrak{p}_{1},\dots,\mathfrak{p}_{s}\}$ denote the subset of prime ideals in $k$ corresponding to the finite places in $\mathrm{Ram}(c(q))$ which split in $k(\sqrt{\delta})$. Then the finite places at which $c(q)\otimes_{k}k(\sqrt{\delta})$ is ramified correspond to the $2s$ prime ideals in $k(\sqrt{\delta})$ lying over $\{\mathfrak{p}_{1},\dots,\mathfrak{p}_{s}\}$ [Mac11, §7]. We get the following condition which we will use to distinguish commensurability classes.
Lemma 2.3.
Suppose that $\mathrm{SO}(q)_{\mathcal{O}_{k}}$ and $\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$ are commensurable for admissible quadratic forms $q$ and $q^{\prime}$ of signature $(n,1)$ defined over $k$ with $n$ odd. If $\nu_{\mathfrak{p}}\in\mathrm{Ram}(c(q))$ corresponds to a prime ideal $\mathfrak{p}$ in $k$ which splits in $k(\sqrt{\delta})$, then $\nu_{\mathfrak{p}}\in\mathrm{Ram}(c(q^{\prime}))$.
3. Realizing isometries in quadratic spaces
In this section we gather results of [BF15] giving necessary and sufficient conditions for a quadratic space to admit an isometry with a given characteristic polynomial.
Definition 3.1.
A monic polynomial $F\in k[x]$is said to be $\epsilon$-symmetric for some $\epsilon=\pm 1$ if $F(x)=\epsilon x^{\deg(F)}F(x^{-1})$. If $\epsilon=1$, we say that $F$ is symmetric.
It is known that minimal polynomials and characteristic polynomials of isometries of quadratic spaces are $\epsilon$-symmetric, with $\epsilon$ equal to the constant term of the polynomial (see [BF15, Proposition 11]). If $F\in k[x]$ is a monic polynomial such that $F(0)\neq 0$, set
$F^{*}(x)=\frac{1}{F(0)}x^{\deg(F)}F(x^{-1}).$
Note that $F^{*}$ is also monic, $F^{*}(0)\neq 0,$ and $F^{**}=F.$
Definition 3.2.
Let $F\in K[x]$ be a monic, $\epsilon$-symmetric polynomial. We say that $F$ is of
  1. type 0:
    if $F$ is a product of powers of $x-1$ and $x+1$.
  2. type 1:
    if $F$ is a product of powers of monic, symmetric, irreducible polynomials in $k[x]$ of even degree.
  3. type 2:
    if $F$ is a product of polynomials of the form $gg^{*}$, where $g\in k[x]$ is monic, irreducible, and $g\neq g^{*}$.
Proposition 3.3.
[BF15, Prop 1.3] Every monic, $\epsilon$-symmetric polynomial $F\in k[x]$ is a product of polynomials of type 0, 1, and 2.
A monic $\epsilon$-symmetric polynomial is hyperbolic if all its factors of type 0 and 1 are of the form $f^{e}$ with $e$ even.
Let $q$ be a quadratic form over $k$ and $F\in k[x]$ be a symmetric polynomial with $\deg(F)=\dim(q)$. For every real place $\nu$ of $k$, let $(r_{\nu},s_{\nu})$ denote the signature of $q_{\nu}$, and let $\sigma_{\nu}$ be the number of pairs of roots of $F\in k_{\nu}[x]$ that are not on the unit circle. Then $F$ and $q$ are said to satisfy the signature condition if for every real place $\nu$ of $k$
$r_{\nu}\geq\sigma_{\nu},\>s_{\nu}\geq\sigma_{\nu},\>\text{and }(r_{\nu},s_{\nu})\equiv(\sigma_{\nu},\sigma_{\nu})\bmod 2.$
If the dimension of $q$ is even and $F$ is as above with $\deg(F)=\dim(q)=2m$, let $S$ be the set of places of $k$ at which the Hasse invariant of $q$ is different from the Hasse invariant of the $2m$-dimensional space associated with the quadratic form $\langle 1,\dots,1,-1,\dots,-1\rangle$ of signature $(m,m)$. By the classification of quaternion algebras, this set $S$ is finite. Then $F$ and $q$ are said to satisfy the hyperbolicity condition if whenever $\nu\in S$, then $F\in k_{\nu}[x]$ is not hyperbolic. (Note that this definition varies slightly from the definition in [BF15] but is implied by [BF15, Corollary 9.3]).
For later use, we note here that the Hasse invariant of the $2m$-dimensional space associated, with the quadratic form $\rho^{m}_{m}=\langle 1,\dots,1,-1,\dots,-1\rangle$ of signature $(m,m)$, is given by
(2) $s(\rho^{m}_{m})=\begin{cases}M_{2}(k)&\text{if }m\equiv 0,1\bmod 4\\ \left(\tfrac{-1,-1}{k}\right)&\text{if }m\equiv 2,3\bmod 4.\end{cases}$
Theorem 3.4.
[BF15, Corollary 9.3] Suppose that $k$ is a global field, $F\in k[x]$ is an irreducible polynomial of type 1, and $q$ is a quadratic form with $\dim(q)=\deg(F)$. The quadratic space $(V,q)$ has an isometry with minimal polynomial $F$ if and only if the following conditions are satisfied:
  1. (i)
    $\det(q)\equiv F(1)F(-1)$ in $k^{\times}/k^{\times 2}$,
  2. (ii)
    the signature condition holds,
  3. (iii)
    the hyperbolicity condition holds.
Bayer-Fluckiger also established in [BF15] conditions for a polynomial of mixed type to be the characteristic polynomial in a quadratic space $(V,q)$. We state the necessary results therein, applied to our cases of interest. We will consider polynomials of the form $F(x)=f(x)\cdot f_{0}(x)$ with $f$ an irreducible polynomial of type 1 and $f_{0}$ a polynomial of type 0.
Proposition 3.5.
[BF15, Proposition 11.11] Suppose that $k$ is a global field, $F(x)=f(x)\cdot f_{0}(x)\in k[x]$ is the product of an irreducible polynomial $f$ of type 1 and a degree 1 polynomial $f_{0}$ of type 0. Let $q$ be a quadratic form of odd dimension with $\dim(q)=\deg(F)=\deg(f)+1$. The quadratic space $(V,q)$ has an isometry with minimal polynomial $F$ if and only if $q\simeq\langle d_{0}\rangle\oplus q^{\prime}$ where $d_{0}=\det(q)f(1)f(-1)$ and $(V^{\prime},q^{\prime})$ has an isometry with characteristic polynomial $f$.
The Witt index of a quadratic form $q$ is the largest integer $m$ such that $q$ has a decomposition with $\rho^{m}_{m}$ (as in Equation 2) as a summand.
Proposition 3.6.
[BF15, Proposition 11.9] Suppose that $k$ is a global field, $F(x)=f(x)\cdot f_{0}(x)\in k[x]$ is the product of an irreducible polynomial $f$ of type 1 and a degree 2 polynomial $f_{0}$ of type 0. Let $q$ be a quadratic form of even rank with $\mathrm{rank}(q)=\deg(F)=\deg(f)+2$. The quadratic space $(V,q)$ has an isometry with minimal polynomial $F$ if and only if the following conditions are satisfied:
  1. (i)
    the signature condition holds
  2. (ii)
    if $\nu$ is a finite place of $k$ where $f\in k_{\nu}[x]$ is hyperbolic, then the Witt index of $q$ defined over $k_{\nu}$ is $\geq\frac{1}{2}\deg(f)$.
Proposition 3.7.
[BF15, Proposition 11.2] Suppose that $k$ is a global field, $F(x)=f(x)\cdot f_{0}(x)\in k[x]$ is the product of an irreducible polynomial $f$ of type 1 and a polynomial $f_{0}$ of type 0 having $\deg(f_{0})\geq 3$. Let $q$ be a quadratic form of odd dimension with $\dim(q)=\deg(F)$. The quadratic space $(V,q)$ has an isometry with minimal polynomial $F$ if and only if the signature condition is satisfied.
The applications of results by Bayer-Fluckiger above give us conditions for the existence of isometries in $\mathrm{O}(q)_{\mathcal{O}_{k}}$ realizing a desired characteristic polynomial. The following lemma realizes such isometries in the integer points, up to $k$-equivalence.
Lemma 3.8.
Let $T$ be an isometry in $\mathrm{O}(q)_{k}$ where $q$ is an admissible quadratic form over $k$. Suppose also that its characteristic polynomial $p_{T}(x)$ has coefficients in the ring of integer $\mathcal{O}_{k}$. Then a conjugate of $T$ is contained in some $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ where $q^{\prime}\simeq q$.
Proof.
Let $r$ denote the rank of $q$, which is also the degree of $p_{T}\in\mathcal{O}_{k}[x]$. Denote the coefficients of $p_{T}$ by $c_{0}=1,c_{1},\dots c_{r-1},c_{r}=1$. The matrix powers $T,T^{2},\dots,T^{r-1}$ all have entries in $k$. Choose $\mathfrak{m}$ to be a common multiple of the primes appearing in the denominators of the entries in these $r-1$ matrices. By the Cayley-Hamilton theorem $T$ satisfies its characteristic polynomial, so for any $a\in\mathbb{Z}_{\geq 0}$
$T^{a+r}=-\left(T^{a}+c_{1}T^{a+1}+\cdots+c_{r-1}T^{a+r-1}\right).$
The subgroup of $k^{r}$ generated by the images $T^{a}(\mathcal{O}_{k}^{r})$ for all $a\geq 0$ lies between the subgroups $\mathcal{O}_{k}^{r}$, and $(\frac{1}{\mathfrak{m}}\mathcal{O}_{k})^{r}$ and is thus itself an $\mathcal{O}_{k}$-lattice of rank $r$. We can therefore choose a change of basis matrix $g\in\mathrm{GL}(r)_{k}$ such that $g^{-1}Tg$ has entries in $\mathcal{O}_{k}$.
If $S_{q}$ is the symmetric matrix associated to the quadratic form $q$, let $q^{\prime}$ be the quadratic form associated to the symmetric matrix $S_{q^{\prime}}=g^{t}S_{q}g$. Then $g^{-1}Tg\in\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ and since $q\simeq q^{\prime}$, $\mathrm{O}(q)_{\mathcal{O}_{k}}$ and $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ are commensurable. ∎
4. Infinitely many commensurability classes
Let $\lambda$ be a Salem number and $f_{\lambda}\in\mathbb{Z}[x]$ its minimal polynomial of degree $2m$. Over the splitting field (or over $\mathbb{C}$), $f_{\lambda}$ factors as
(3) $f_{\lambda}(x)=(x-\lambda)(x-\lambda^{-1})(x-e^{i\theta_{1}})(x-e^{-i\theta_{1}})\cdots(x-e^{i\theta_{m-1}})(x-e^{-i\theta_{m-1}}),$
which equals
(4) $f_{\lambda}(x)=(x^{2}-(\lambda+\lambda^{-1})x+1)(x^{2}-2\cos\theta_{1}x+1)\cdots(x^{2}-2\cos\theta_{m-1}x+1).$
Over the $\mathbb{R}$, $f_{\lambda}$ factors as
(5) $f_{\lambda}(x)=(x-\lambda)(x-\lambda^{-1})(x^{2}-2\cos\theta_{1}x+1)\cdots(x^{2}-2\cos\theta_{m-1}x+1).$
In what follows, we fix the following notation.
  • $n\geq 2$.
  • $\lambda$ a Salem number with minimal polynomial $f_{\lambda}\in\mathbb{Z}[x]$ over $\mathbb{Q}$.
  • $k$ is a number field with $k\subseteq\mathbb{Q}(\lambda+\lambda^{-1})$.
  • $f(x)\in k[x]$ the minimal polynomial of $\lambda$ over $k$.
  • The number field $K=\mathbb{Q}(\lambda)=k(\lambda)\cong k[x]/(f)$ with splitting field $K^{\prime}$.
  • $g(x)\in k[x]$ the minimal polynomial of $\lambda+\lambda^{-1}$ over $k$.
  • $E=\mathbb{Q}(\lambda+\lambda^{-1})=k(\lambda+\lambda^{-1})\cong k[x]/(g)$ with splitting field $E^{\prime}$.
  • $D=f(1)f(-1)$.
  • $\delta\equiv(-1)^{n(n+1)/2}D$ as an element in $k^{\times}/k^{\times 2}$.
  • $h(x)=x^{2}-\delta\in k[x]$ the minimal polynomial of $\delta$ over $k$.
  • $H=k(\sqrt{\delta})\cong k[x]/(h)$.
  • $M=E^{\prime}H$ the compositum of $E^{\prime}$ and $H$.
  • $L=K^{\prime}M$ the compositum of $K^{\prime}$ and $M$.
Note that the fields $H$, $M$, and $L$ are Galois extensions of $k$.
We will prove the Main Theorem in several cases according to the difference $n+1-\deg(f)$ as well as the parity of $n$. The case that $n+1-\deg(f)\geq 3$ is the easiest. Before presenting it, we need to guarantee some conditions over $D$.
Lemma 4.1.
$D<0$ and if $\sigma$ is a non-identity embedding $\sigma:k\hookrightarrow\mathbb{R}$, then $\sigma(D)>0$.
Proof.
By [ERT19, Lemmas 6.1 and 6.2], $2m=\deg(\lambda)=\deg_{k}(\lambda)\cdot[k:\mathbb{Q}]$, $\deg_{k}(\lambda)=2m_{k}$ for some integer $m_{k}|m$. We have that $f|f_{\lambda}$ and the roots of $f$ are a subset of the reciprocal pairs of roots of $f_{\lambda}$. Over $\mathbb{R}$, $f_{\lambda}$ factors as in Equation 5 and then $f$ factors a subset of $m_{k}$ of these, including $f_{0}=\left(x^{2}-(\lambda+\lambda^{-1})x+1\right)$. All other factors of $f_{\lambda}$ have the form $f_{i}=(x^{2}-2\cos\theta_{i}x+1)$ for $i=1,\dots,m-1$. Clearly $f_{0}(1)f_{0}(-1)<0$ and for $i\not=0$ $f_{i}(1)f_{i}(-1)>0$. A non-identity embedding will send all the factors of $f$ to factors in $f_{\lambda}/f$. So $f(1)f(-1)<0$ and $\sigma\left(f(1)f(-1)\right)>0$. ∎
Proposition 4.2.
If $n+1-\deg(f)\geq 3$ then every commensurability class of arithmetic lattice of simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ realizes $\lambda$.
Proof.
Consider an arbitrary admissible quadratic from $q$ over $k$ with signature $(n,1)$. Let $F(x)=f(x)\cdot(x+1)^{n+1-\deg(f)}$. By Proposition 3.7, there is an isometry $T\in\mathrm{O}(q)_{k}$ with characteristic polynomial $p_{T}=F$ if and only if the signature condition holds. That the signature condition holds follows from the proof of Lemma 4.1.
By Lemma 3.8, a conjugate $\gamma$ of $T$ is contained in $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ for some $q^{\prime}\simeq q$. Note that $\det(\gamma)=\det(T)=F(0)=1$, so $\gamma\in\Gamma=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$. Thus $\lambda$ is realized in the commensurability class of $\mathrm{SO}(q)_{\mathcal{O}_{k}}$. ∎
The cases with $n+1-\deg(f)\leq 2$ will be treated depending on the parity of $n$.
4.1. Some algebraic lemmas
We will need several lemmas.
Lemma 4.3.
[BF15, Lemma 9.4] Let $\nu$ be a place of $k$. Then $f\in k_{\nu}[x]$ is hyperbolic if and only if for any place $\nu^{\prime}$ of $E$ lying above $\nu$, $\nu^{\prime}$ splits in $K$.
Let $\Sigma^{ns}(K)$ denote the set of prime ideals $\mathfrak{p}$ in $k$ for which there exists a prime ideal Weierstrass-$\wp$ in $E$ lying over $\mathfrak{p}$ which remains inert in $K/E$.
Whenever $M$ is a field extension of $k$, denote by $\mathrm{Spl}(M/k)$ the set of prime ideals in $k$ that split completely in $M$. Whenever $S$ is any set of prime ideals in $k$ (usually finite), $\mathrm{Spl}_{S}(M/k):=\mathrm{Spl}(M/k)-S$. The following lemma is a well-known extension of a theorem originally due to Bauer. We include a proof using the Chebotarev density theorem [Tsc26].
Lemma 4.4.
Let $L$ and $M$ be Galois extensions of $k$. Suppose $S$ is a finite set of prime ideals in $k$. Then $L\subseteq M$ if and only if $\mathrm{Spl}_{S}(M/k)\subseteq\mathrm{Spl}_{S}(L/k)$.
Proof.
One direction is immediate. To prove the other direction, note first that
$\mathrm{Spl}_{S}(LM/k)=\mathrm{Spl}_{S}(L/k)\cap\mathrm{Spl}_{S}(M/k).$
[…]
Lemma 4.5.
$K\not\subseteq M$ except possibly in the case that $\deg_{k}(\lambda)=2$ and $\delta=-D$.
Proof.
We first note that $E^{\prime}$ is a real field (in fact, totally real) since the conjugates over $\mathbb{Q}$ of $\lambda+\lambda^{-1}$ are $2\cos\theta_{j}$ for $11\leq j\leq m-$ 1, and thus the conjugates over $k$ are some subset of those. By construction, $\lambda\not\in E^{\prime}$. Then $E^{\prime}(\lambda)$ is quadratic over $E^{\prime}$.
If $\delta$ is a square in $E^{\prime}$, then $M=E^{\prime}$ and it is clear that $K\not\subseteq M$.
Now assume $\delta$ is not a square in $E^{\prime}$. Then $M=E^{\prime}(\sqrt{\delta})$ is quadratic over $E^{\prime}$, as is $E^{\prime}(\lambda)$. Assume that $K\subseteq M$, then $E^{\prime}\subseteq E^{\prime}(\lambda)\subseteq M$, which implies $E^{\prime}(\lambda)=M$.
Suppose that $\deg_{k}(\lambda)=2$ and $\delta=D<0$. This implies that $k=E=E^{\prime}$ and $f(x)=x^{2}-(\lambda+\lambda^{-1})x+1$. Then $K$ and $M$ are both of degree 2 over $E$, one of which is real and the other of which is not real, which is a contradiction.
Now suppose that $\deg_{k}(\lambda)\geq 4$. Then $\lambda$ has non-real conjugates over $k$. By construction $M$ is Galois over $k$. But $E^{\prime}(\lambda)$ is not Galois over $k$ since it is real and hence does not contain any root of $f$ other than $\lambda$ and $\lambda^{-1}$. Again, this is a contradiction. ∎
4.2. Completing the proof
We now complete the proof of the Main Theorem, which will follow from Propositions 4.2, 4.6, 4.7, 4.8, 4.9.
We treat first the case of $n=2$, which gives a similar result to [MR03, Corollary 12.2.10].
Proposition 4.6.
Let $n=2$. If $\deg(f)=2$ then there exist infinitely many commensurability classes of arithmetic lattice of simplest type in $\mathrm{Isom}(\mathbb{H}^{2})$ defined over $k$ which realize $\lambda$.
Proof.
Under these hypothesis, $K=L$, $E=k$, $M=H$, and $\delta=-D$. If $K=M$ then $\Sigma^{ns}(K)-\mathrm{Spl}(H/k)=\Sigma^{ns}(K)$. Otherwise, $K$ and $M$ are distinct Galois extensions of $k$, both of degree two. It follows from Lemma 4.4 and the Chebotarev density theorem that in either case we have #$\#\left(\Sigma^{ns}(K)-\mathrm{Spl}(M/k)\right)=\infty$.
Choose a finite subset of even cardinality $\mathcal{A}$ satisfying
$\mathcal{A}\subseteq\left(\Sigma^{ns}(K)-\mathrm{Spl}(M/k)\right).$
Whenever $\nu\in\mathcal{A}$, $\nu\not\in\mathrm{Spl}(k(\sqrt{\delta})/k)$ since $M=k(\sqrt{\delta})$. This means $\delta$ is a non-square at $k_{\nu}$. Therefore, by [O’M00, Theorem 71:19], there exists $a\in k^{\times}$ such that the quaternion algebra $\left(\frac{a,\delta}{k}\right)$ is ramified exactly at $\mathcal{A}$. Define the admissible quadratic form
$q=\langle-\delta a,a,1\rangle$
which has Hasse invariant $\left(\frac{a,-a\delta}{k}\right)\cong\left(\frac{a,\delta}{k}\right)$.
Let $F(x)=f(x)\cdot(x+1)$. By Proposition 3.5, since $1\equiv D^{2}$ in $k^{\times}/k^{\times 2}$ and $q\simeq\langle 1\rangle\oplus\langle-\delta a,a\rangle$, there is an isometry in $\mathrm{O}(q)_{\mathcal{O}_{k}}$ with characteristic polynomial $f(x)\cdot(x+1)$ if and only if there is an isometry $T\in\mathrm{O}(\langle-\delta a,a\rangle)$ with characteristic polynomial $f(x)$ if and only if the conditions in Theorem 3.4 are satisfied.
Condition $(i)$ is satisfied since
$\det(\langle-\delta a,a\rangle)=a^{2}D\equiv D\equiv f(1)f(-1)\text{ in }k^{\times}/k^{\times 2}.$
Condition $(ii)$ is satisfied by the proof of Lemma 4.1.
Finally, condition $(iii)$ is satisfied by our choice of $\mathcal{A}$ and Lemma 4.3. Indeed, if $f$ is hyperbolic in $k_{\nu}[x]$, then $\nu\not\in\Sigma^{ns}(K)$ by Lemma 4.3 and by Equation 2 the set $S$ from the definition of hyperbolicity is equal to $\mathcal{A}\subseteq\Sigma^{ns}(K)$. Thus, $f$ is not hyperbolic at any place in $S$.
By Lemma 3.8, a conjugate $\gamma$ of $T$ is contained in $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ for some $q^{\prime}\simeq q$. Note that $\det(\gamma)=\det(T)=F(0)=1$, so $\gamma\in\Gamma=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$. Varying over different choices of $\mathcal{A}$, we get infinitely many quadratic forms $q$’s with pairwise distinct Hasse invariant, and hence pairwise distinct Witt invariant. By Proposition 2.2, this gives infinitely many commensurability classes. ∎
Proposition 4.7.
Let $n\geq 4$ be even. If $n+1-\deg(f)=1$ then there exist infinitely many commensurability classes of arithmetic lattice of simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ which realize $\lambda$.
Proof.
We claim first that #$\#\Sigma^{ns}(K)=\infty$.
The extension $K/E$ is a Galois extension of degree 2. By the Chebotarev density theorem, the set of prime ideals Weierstrass-$\wp$ in $E$ which remain inert in $K$ has density $1/2$. At most $[E:k]$-many such Weierstrass-$\wp$ can lie over the same prime ideal $\mathfrak{p}=\wp\cap\mathcal{O}_{k}$ of $k$. Thus $\Sigma^{ns}(K)$ is an infinite set.
Denote the standard positive definite quadratic form of rank $r$ and signature $(r,0)$ by $Q_{r}=\langle 1,\dots,1\rangle$. Choose a finite subset of even cardinality $\mathcal{A}$ satisfying
$\mathcal{A}\subseteq\Sigma^{ns}(K),\text{ if }\frac{n}{2}\equiv 0,1\bmod 4$
and
$\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right)\subseteq\mathcal{A}\subseteq\Sigma^{ns}(K)\cup\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right),\text{ if }\frac{n}{2}\equiv 2,3\bmod 4.$
Let $B_{\mathcal{A}}$ be the quaternion algebra with ramification set $\mathcal{A}$, which only contains finite places. Choose $a,b\in k^{\times}$ so that
$\left(\tfrac{a,b}{k}\right)=B_{\mathcal{A}}\cdot\left(\tfrac{-1,-D}{k}\right).$
Note that $\left(\tfrac{a,b}{k}\right)$ is ramified at all real places except $\nu_{0}$. By Lemma 4.1, at $\nu_{0}$ we have $D<0$ and at least one of $a$ or $b$ is positive. At any other real place $\nu$, $D>0$, $a<0$, and $b<0$. Define the admissible quadratic form
$q=\langle-Da,-Db,Dab\rangle\oplus Q_{n-2}$
which has Hasse invariant
$s(q)=s\left(\langle-Da,-Db,Dab\rangle\right)=\left(\tfrac{a,b}{k}\right)\cdot\left(\tfrac{-1,-D}{k}\right)=B_{\mathcal{A}}.$
Note that $q$ has the correct signatures and is admissible.
Let $F(x)=f(x)\cdot(x+1)$. By Proposition 3.5, since $1\equiv D^{2}$ in $k^{\times}/k^{\times 2}$, there is an isometry in $\mathrm{O}(q)_{\mathcal{O}_{k}}$ with characteristic polynomial $F$ if and only if there is an isometry in $\mathrm{O}(\langle-Da,-Db,Dab\rangle\oplus Q_{n-3})$ with characteristic polynomial $f(x)$ if and only if the conditions in Theorem 3.4 are satisfied.
Condition $(i)$ is satisfied since
$\det(\langle-Da,-Db,Dab\rangle)\equiv D\equiv f(1)f(-1)\text{ in }k^{\times}/k^{\times 2}.$
Condition $(ii)$ is satisfied by the proof of Lemma 4.1.
We claim condition $(iii)$ is satisfied by our choice of $\mathcal{A}$. Indeed, if $\frac{n}{2}\equiv 0,1\bmod 4$ and $f$ is hyperbolic in $k_{\nu}[x]$, then $\nu\not\in\Sigma^{ns}(K)$ by Lemma 4.3 and by eq. 2 the set $S$ from the definition of hyperbolicity is equal to $\mathcal{A}\subseteq\Sigma^{ns}(K)$.
Instead, if $\frac{n}{2}\equiv 2,3\bmod 4$, the set $S$ is equal to $\Omega_{\infty}\cup\left(\mathcal{A}-\mathrm{Ram}_{fin}\left(\frac{-1,-1}{k}\right)\right)$. If $\nu$ is a real place then $f$ factors over $k_{\nu}$ into a subset of the factors in eq. 5, with at least one factor of the type $(x^{2}-2\cos\theta_{i}x+1)$ (not repeated), since $\deg_{k}(f)\geq 4$. So $f$ is not hyperbolic at any real place. If $\nu\in S$ is a finite place, then $\nu\in\Sigma^{ns}(K)$, so $f$ is not hyperbolic at $\nu$. In either case, $f$ is not hyperbolic at any place in $S$ and thus, condition $(iii)$ is satisfied.
By Lemma 3.8, a conjugate $\gamma$ of $T$ is contained in $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ for some $q^{\prime}\simeq q$. Note that $\det(\gamma)=\det(T)=F(0)=1$, so $\gamma\in\Gamma=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$. Varying over different choices of $\mathcal{A}$, we get infinitely many quadratic forms $q$’s with pairwise distinct Hasse invariant, and hence pairwise distinct Witt invariant. By Proposition 2.2, this gives infinitely many commensurability classes. ∎
Proposition 4.8.
Let $n\geq 3$ be odd. If $n+1-\deg(f)=2$ then there exist infinitely many commensurability classes of arithmetic lattices of simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ which realize $\lambda$.
Proof.
As in the beginning of the proof of Proposition 4.6, it follows from Lemmas 4.4 and 4.5 that
#$\#\left(\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)\right)=\infty.$
Denote the standard positive definite quadratic form of rank $r$ signature $(r,0)$ by $Q_{r}=\langle 1,\dots,1\rangle$. Choose $\mathcal{A}$ to be a finite subset of even cardinality such that if $\frac{n-1}{2}\equiv 0,1\bmod 4$, then
$\mathcal{A}\subseteq\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)$
and if $\frac{n-1}{2}\equiv 2,3\bmod 4$, then
$\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right)\subseteq\mathcal{A}\subseteq\left(\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)\right)\cup\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right).$
Let $B_{\mathcal{A}}$ be the quaternion algebra with ramification set $\mathcal{A}$, which only contains finite places. Choose $a,b\in k^{\times}$ so that
$\left(\tfrac{a,b}{k}\right)=B_{\mathcal{A}}\cdot\left(\tfrac{-1,-D}{k}\right).$
Note that $\left(\tfrac{a,b}{k}\right)$ is ramified at all real places except $\nu_{0}$. By Lemma 4.1, at $\nu_{0}$ we have $D<0$ and at least one of $a$ or $b$ is positive. At any other real place $\nu$, $D>0$, $a<0$, and $b<0$. Define the admissible quadratic form
$q=\langle-Da,-Db,Dab\rangle\oplus Q_{n-2}$
which has Hasse invariant
$s(q)=s\left(\langle-Da,-Db,Dab\rangle\right)=\left(\tfrac{a,b}{k}\right)\cdot\left(\tfrac{-1,-D}{k}\right)=B_{\mathcal{A}}.$
Note that $q$ has the correct signatures and is admissible.
Let $F(x)=f(x)\cdot(x+1)^{2}$. By Proposition 3.6, there is an isometry in $\mathrm{O}(q)_{\mathcal{O}_{k}}$ with characteristic polynomial $F(x)$ if and only if the conditions $(i)-(ii)$ are satisfied. Condition (i) is satisfied by the proof of Lemma 4.1.
If $\frac{n-1}{2}\equiv 0,1\bmod 4$ and $f$ is hyperbolic in $k_{\nu}[x]$, then $\nu\not\in\Sigma^{ns}(K)$ by Lemma 4.3 and by eq. 2 the set $S$ from the definition of hyperbolicity is equal to $\mathcal{A}\subseteq\Sigma^{ns}(K)$.
If $\frac{n-1}{2}\equiv 2,3\bmod 4$ the set $S$ is equal to $\Omega_{\infty}\cup\left(\mathcal{A}-\mathrm{Ram}\left(\frac{-1,-1}{k}\right)\right)$. As in the proof of Proposition 4.7 $f$ is not hyperbolic at any real place. If $\nu\in\mathcal{A}-\mathrm{Ram}\left(\frac{-1,-1}{k}\right)\subseteq\Sigma^{ns}(K)$ then by Lemma 4.3, $f$ is not hyperbolic at $\nu$.
In either case, if $f$ is hyperbolic at $k_{\nu}$, then $\nu\not\in S$. Also, $f$ hyperbolic at $\nu$ implies that there exists a polynomial $p\in k_{\nu}[x]$ with $\deg p=\frac{n-1}{2}$ such that $f=pp^{*}$. Then
$D\equiv f(1)f(-1)\equiv p(1)^{2}\cdot(-1)^{(n-1)/2}\cdot p(-1)^{2}\text{ in }k^{\times}/k^{\times 2}.$
Therefore the form $\langle-Da,-Db,Dab\rangle\oplus Q_{n-2}$ over $k_{\nu}$ has the same Hasse-invariant, determinant (up to squares), and dimension as the form $\rho^{(n-1)/2}_{(n-1)/2}$ from eq. 2, which has Witt index $(n-1)/2$. Thus, over $k_{\nu}$, the quadratic form $q$ has Witt index $\geq(n-1)/2=\frac{1}{2}\deg(f)$, and condition $(ii)$ is satisfied.
By Lemma 3.8, a conjugate $\gamma$ of $T$ is contained in $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ for some $q^{\prime}\simeq q$. Note that $\det(\gamma)=\det(T)=F(0)=1$, so $\gamma\in\Gamma=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$. Varying over different choices of $\mathcal{A}$, we get infinitely many quadratic forms $q$’s with pairwise distinct Hasse and Witt invariants. Furthermore, if $\mathfrak{p}\in\mathcal{A}$ but $\mathfrak{p}\not\in\mathcal{A}^{\prime}$ for two such $q$ and $q^{\prime}$ constructed as above, then since $\mathfrak{p}\in\mathrm{Spl}(H/k)$, by Lemma 2.3 $\mathrm{SO}(q)_{\mathcal{O}_{k}}$ and $\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$ are incommensurable.
Again, this gives infinitely many commensurability classes. ∎
Proposition 4.9.
Let $n\geq 3$ be odd. If $n+1=\deg(f)$ then there exist infinitely many commensurability classes of arithmetic lattice of simplest type in $\mathrm{Isom}(\mathbb{H}^{n})$ defined over $k$ which realize $\lambda$.
Proof.
It follows from Lemmas 4.4 and 4.5 that
#$\#\left(\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)\right)=\infty.$
Denote the standard positive definite quadratic form of rank $r$ and signature $(r,0)$ by $Q_{r}=\langle 1,\dots,1\rangle$. Choose $\mathcal{A}$ to be a finite subset of even cardinality such that if $\frac{n+1}{2}\equiv 0,1\bmod 4$, then
$\mathcal{A}\subseteq\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)$
and if $\frac{n+1}{2}\equiv 2,3\bmod 4$, then
$\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right)\subseteq\mathcal{A}\subseteq\left(\Sigma^{ns}(K)\cap\mathrm{Spl}(H/k)\right)\cup\mathrm{Ram}_{fin}\left(\tfrac{-1,-1}{k}\right).$
Define the admissible quadratic form
$q=\langle-Da,-Db,Dab\rangle\oplus Q_{n-2}$
which has Hasse invariant
$s(q)=s\left(\langle-Da,-Db,Dab\rangle\right)=\left(\tfrac{a,b}{k}\right)\cdot\left(\tfrac{-1,-D}{k}\right)=B_{\mathcal{A}}.$
Note that $q$ has the correct signatures and is admissible.
By Theorem 3.4, there is an isometry in $\mathrm{O}(q)_{\mathcal{O}_{k}}$ with characteristic polynomial $f(x)$ if and only if the conditions $(i)-(iii)$ are satisfied. Condition $(i)$ is satisfied since
$\det(\langle-Da,-Db,Dab\rangle)\equiv D\equiv f(1)f(-1)\text{ in }k^{\times}/k^{\times 2}.$
Condition $(ii)$ is satisfied by the proof of Lemma 4.1.
If $\frac{n+1}{2}\equiv 0,1\bmod 4$ and $f$ is hyperbolic in $k_{\nu}[x]$, then $\nu\not\in\Sigma^{ns}(K)$ by Lemma 4.3 and by eq. 2 the set $S$ from the definition of hyperbolicity is equal to $\mathcal{A}\subseteq\Sigma^{ns}(K)$.
If $\frac{n+1}{2}\equiv 2,3\bmod 4$, the set $S$ is equal to $\Omega_{\infty}\cup\left(\mathcal{A}-\mathrm{Ram}\left(\frac{-1,-1}{k}\right)\right)$. As before, $f$ is not hyperbolic at any real place. Similarly, by Lemma 4.3, $f$ is not hyperbolic at any $\nu\in\mathcal{A}-\mathrm{Ram}\left(\frac{-1,-1}{k}\right)\subseteq\Sigma^{ns}(K)$. In either case, if $f$ is hyperbolic at $k_{\nu}$, then $\nu\not\in S$ and condition $(iii)$ is satisfied.
By Lemma 3.8, a conjugate $\gamma$ of $T$ is contained in $\mathrm{O}(q^{\prime})_{\mathcal{O}_{k}}$ for some $q^{\prime}\simeq q$. Note that $\det(\gamma)=\det(T)=F(0)=1$, so $\gamma\in\Gamma=\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$. Varying over different choices of $\mathcal{A}$, we get infinitely many quadratic forms $q$’s with pairwise distinct Hasse and Witt invariants. Furthermore, if $\mathfrak{p}\in\mathcal{A}$ but $\mathfrak{p}\not\in\mathcal{A}^{\prime}$ for two such $q$ and $q^{\prime}$ constructed as above, then since $\mathfrak{p}\in\mathrm{Spl}(H/k)$, by Lemma 2.3 $\mathrm{SO}(q)_{\mathcal{O}_{k}}$ and $\mathrm{SO}(q^{\prime})_{\mathcal{O}_{k}}$ are incommensurable. Again, this gives infinitely many commensurability classes. ∎